OpenStudy (anonymous):

What is the equation of the line that passes through the points (−3, 2) and (−5, 8)? 3x − y = −11 3x + y = − 7 3x + y = 5 x + 3y = 3

OpenStudy (compassionate):

First you want to find the slope. Do you know the slope formula, Chrstine?

OpenStudy (anonymous):

yes slope=y2-y1/x2-x1

OpenStudy (anonymous):

slope=-3

OpenStudy (compassionate):

Alright. Now take that -3 slope and use the formula y - y1 = m(x - x1)

OpenStudy (anonymous):

y-2=-3(x-(-3)) y-2=-3x-9 +2 +2 y=-3x-7 is that correct?

OpenStudy (compassionate):

Now put it in Standard notation.

OpenStudy (anonymous):

y=-3x-7 +3x +3x 3x+y=-7 is that the answer?

OpenStudy (anonymous):

THANK YOU!!!!